# Are You Doubtful? Oh, It Might Be Difficult Then! Exploring the Use of Model Uncertainty for Question Difficulty Estimation

### v1

@@ Appendix A Model Correctness and Uncertainty
Table 4 presents the performance of each model on the two question sets, as well as the relation between model certainty and model correctness. In line with the results of Plaut et al. 2024, it is clear that both tested metrics correlate well with model correctness: For all models, on average, the mean certainty is significantly higher for the correctly answered question items. This suggests that the two metrics indeed capture an aspect of model certainty.
Dataset Model Overall Correctness Mean Probability
1st Token Choice Order
Biopsychology phi3_5-chat 0.302 0.448 / 0.232 0.532 / 0.240
Llama3_2-3b-chat 0.707 0.733 / 0.247 0.853 / 0.196
Qwen2_5-3b-chat 0.799 0.864 / 0.200 0.883 / 0.192
Llama3_1-8b-chat 0.824 0.683 / 0.266 0.860 / 0.236
Qwen2_5-14b-chat 0.902 0.972 / 0.136 0.975 / 0.115
Qwen2_5-32b-chat 0.932 0.968 / 0.154 0.978 / 0.121
Yi-34b-chat 0.852 0.878 / 0.234 0.907 / 0.215
Qwen2_5-72b-chat 0.937 0.964 / 0.175 0.984 / 0.139
Llama3_1-70b-chat 0.933 0.946 / 0.209 0.980 / 0.160
BEA phi3_5-chat 0.193 0.368 / 0.158 0.426 / 0.152
Llama3_2-3b-chat 0.645 0.607 / 0.183 0.782 / 0.153
Qwen2_5-3b-chat 0.510 0.676 / 0.148 0.752 / 0.123
Llama3_1-8b-chat 0.654 0.441 / 0.194 0.758 / 0.169
Qwen2_5-14b-chat 0.741 0.905 / 0.163 0.915 / 0.150
Qwen2_5-32b-chat 0.805 0.912 / 0.170 0.936 / 0.151
Yi-34b-chat 0.651 0.748 / 0.175 0.802 / 0.155
Qwen2_5-72b-chat 0.855 0.916 / 0.195 0.953 / 0.169
Llama3_1-70b-chat 0.892 0.845 / 0.181 0.945 / 0.144
@@
Table 4: Model correctness and answer probability in terms of Mean 1st Token and Choice Order Probability in the Biopsychology and BEA question sets. “Overall Correctness" indicates the proportion of correctly answered questions, while the probabilities in green and red indicate the mean model certainty for correctly and incorrectly answered questions respectively. As can be seen, on average, model certainty is significantly higher for questions that are answered correctly.
Appendix B LLM-as-a-Judge Approach
In this section we briefly explore the possibility of using a strong LLM, Llama3.1-70B-Chat, to tackle the task of MCQ item difficulty estimation. Specifically, for each of the two datasets we instruct the LLM to predict the number of students correctly answering the given question, using the prompt presented in table 5. This approach results in RMSEs of $0.28810.2881$ and $0.35650.3565$ for the Biopsychology and BEA datasets respectively, which is significantly worse than the mean regressor baseline.
Instruction Prompt
Below is a multiple-choice question. Out of 100 students, how many do you think answered correctly? Answer with a number between 0 and 100 and do not include an explanation or any other text.
Question:
[Question Text]
Number of students (out of 100) answering correctly:
Table 5: Instruction phrasing used for the LLM-as-a-Judge exploratory experiment. [Question Text] is replaced by the item stem followed by the answer choices, each prepended with the corresponding letter A to J.
Appendix C Large Language Models Used
@@
Model Source
phi3_5-chat https://huggingface.co/unsloth/Phi-3.5-mini-instruct-bnb-4bit
Llama3_2-3b-chat https://huggingface.co/unsloth/Llama-3.2-3B-Instruct-bnb-4bit
Qwen2_5-3b-chat https://huggingface.co/unsloth/Qwen2.5-3B-Instruct-bnb-4bit
Llama3_1-8b-chat https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit
Qwen2_5-32b-chat https://huggingface.co/unsloth/Qwen2.5-32B-Instruct-bnb-4bit
Qwen2_5-14b-chat https://huggingface.co/unsloth/Qwen2.5-14B-Instruct-bnb-4bit
Yi-34b-chat https://huggingface.co/unsloth/yi-34b-chat-bnb-4bit
Qwen2_5-72b-chat https://huggingface.co/unsloth/Qwen2.5-72B-Instruct-bnb-4bit
Llama3_1-70b-chat https://huggingface.co/unsloth/Meta-Llama-3.1-70B-Instruct-bnb-4bit
@@
Table 6: LLMs used in the experiments
Appendix D Use of Bio_Clinical Bert Embeddings
@@
Table 7 presents the performance of our setup using the Bio_clinical Bert Embeddings Alsentzer et al. (2019). The results using BERT embeddings Devlin et al. (2019) are repeated to facilitate comparison. In our experiments, using the Bio_clinical Bert Embeddings consistently led to worse performance (higher RMSE) for all experiments.
Biopsychology BEA
Only Text Features
Bio_Clinical BERT 0.1518 0.3094
BERT 0.1498 0.3066
Text and Model Uncertainty Features
Bio_Clinical BERT BERT Bio_Clinical BERT BERT
First Token Probability 0.1396 0.1385 0.2908 0.2854
Choice Order Sensitivity 0.1426 0.1411 0.3035 0.2961
1st Token Probability & Choice Order Sensitivity 0.1392 0.1388 0.2909 0.2846
@@
Table 7: Comparison of results using Bio_Clinical BERT or default BERT embeddings in terms of RMSE on the test set. All results are averaged over ten repetitions, with the standard deviation not exceeding 0.002.
Appendix E Additional SHAP analyses
In this section we present a collection of additional analyses using the Shapley additive explanations Lundberg and Lee (2017).
E.1 Only Using TF-IDF Features using BEA Dataset
Figure 5 presents the effect of the ten most impactful uni/bi-grams on the trained Random Forest’s difficulty prediction. Questions containing the words "physician" and "examination" are generally predicted to be easy, while the other impactful features lead to the Random Forest predicting higher difficulty (lower proportion of students selecting the correct answer).
Figure 5: BEA Dataset. SHAP summary plot showing the contribution of the top ten uni/bi-gram features to the Random Forest’s predictions, highlighting their importance and impact direction. Features are ranked by their average influence, with dots representing individual question items and colour indicating TF-IDF scores. Results averaged over ten repetitions.
E.2 TF-IDF Features and All Model Uncertainty Features
Figure 6 shows the SHAP summary plot when both model uncertainty feature-sets are given to the Random Forest Regressor, along with the text encoded as TF-IDF scores. Here, we observe that overall 1st Token Probabilities are preferred over Choice-Order Probabilities as a proxies for difficulty.
(a) Biopsychology Dataset
(b) BEA Dataset
Figure 6: SHAP summary plots for the Biopsychology and BEA datasets showing the contribution of the top ten features to the Random Forest’s predictions provided TF-IDF scores and all model uncertainty features. Higher First Token Probability and Order Probability metrics indicate greater model certainty. Results averaged over ten repetitions.
Appendix F Using Semantic Similarity Between Answer Choices
While the central goal of our study is to showcase that model uncertainty is a useful signal for MCQ item difficulty, in this section, we show that additional features have the potential to further the performance of our setup. More in detail, we use a choices similarity metric, defined (per MCQ) as the average cosine similarity between each distractor and the correct answer choice, similar to the approach by (Susanti et al., 2020). This is operationalised using the Sentence Transformer library Reimers and Gurevych (2019) with one of two models: "all-MiniLM-L6-v2" [^9] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 (general efficient embeddings) and "S-PubMedBert-MS-MARCO" [^10] https://huggingface.co/pritamdeka/S-PubMedBert-MS-MARCO (medical/health text domain embeddings). The two setups are henceforth refered to as "General Similarity" and "Medical Similarity" respectively.
Table 8 presents the achieved RMSE using text features, model uncertainty and choice similarity in the two datasets.
RMSE
Best Result without Choice Similarity
Biopsychology: TF-IDF & Choice-Order Sensitivity 0.1309
BEA: BERT Embeddings, 1st Token Probability & Choice-Order Sensitivity 0.2846
Biopsychology BEA
Only Choice Similarity Features
General Similarity 0.1895 0.3567
Medical Similarity 0.1883 0.3432
Text, Model Uncertainty and Choice Similarity Features
TF-IDF BERT TF-IDF BERT
1st Token Probability, Choice Order Sensitivity & General Similarity 0.1386 0.1389 0.2850 0.2841
1st Token Probability, Choice Order Sensitivity & Medical Similarity 0.1378 0.1381 0.2856 0.2844
Table 8: Performance of Random Forest Regressor using text, model uncertainty and choice similarity features. Best score per dataset indicated in boldface. All results are averaged over ten repetitions, with the standard deviation not exceeding 0.002.

### v2

@@ Appendix A Model Correctness and Uncertainty
Table 5 presents the performance of each model on the three question sets, as well as the relation between model certainty and model correctness. In line with the results of Plaut et al. [22], it is clear that both tested metrics correlate well with model correctness: On average, the mean certainty for the chosen option is higher for the correctly answered question items. This suggests that the two metrics indeed capture an aspect of model certainty. Lastly, we can see that this trend persists despite quantisation, although quantised models generally exhibit lower performance compared to their default precision counterparts.
@@
Table 5: Model correctness and answer probability in terms of Mean 1st Token and Choice Order Probability in the Biopsychology, USMLE and CMCQRD question sets. “Overall Correctness" indicates the proportion of correctly answered questions, while the probabilities in green and red indicate the mean model certainty for the model’s choice for correctly and incorrectly answered questions respectively. As can be seen, on average, model certainty is higher when questions are answered correctly, especially for larger LLMs.
Default Precision 4-bit Quantisation
Dataset Model Correctness Mean Probability Correctness Mean Probability
1st Token Choice Order 1st Token Choice Order
Biopsychology phi3_5 0.821 0.921 / 0.696 0.940 / 0.743 0.288 0.418 / 0.397 0.494 / 0.455
Llama3_2-3b 0.740 0.747 / 0.541 0.849 / 0.676 0.714 0.735 / 0.509 0.855 / 0.652
Qwen2_5-3b 0.772 0.830 / 0.609 0.856 / 0.641 0.797 0.857 / 0.610 0.874 / 0.642
Llama3_1-8b 0.817 0.641 / 0.445 0.863 / 0.648 0.832 0.677 / 0.455 0.847 / 0.596
Qwen2_5-14b 0.896 0.968 / 0.790 0.972 / 0.823 0.897 0.971 / 0.777 0.973 / 0.800
Qwen2_5-32b 0.934 0.972 / 0.759 0.981 / 0.802 0.933 0.968 / 0.750 0.978 / 0.792
Yi-34b 0.880 0.888 / 0.629 0.917 / 0.681 0.874 0.868 / 0.582 0.896 / 0.615
Llama3_1-70b 0.933 0.938 / 0.649 0.975 / 0.794 0.935 0.945 / 0.690 0.977 / 0.811
Qwen2_5-72b 0.941 0.971 / 0.808 0.981 / 0.875 0.939 0.962 / 0.750 0.982 / 0.815
USMLE phi3_5 0.571 0.781 / 0.612 0.838 / 0.703 0.189 0.319 / 0.306 0.363 / 0.372
Llama3_2-3b 0.634 0.618 / 0.428 0.785 / 0.614 0.658 0.596 / 0.407 0.767 / 0.572
Qwen2_5-3b 0.477 0.670 / 0.563 0.740 / 0.648 0.514 0.663 / 0.529 0.740 / 0.630
Llama3_1-8b 0.627 0.371 / 0.290 0.637 / 0.507 0.679 0.431 / 0.325 0.730 / 0.585
Qwen2_5-14b 0.744 0.897 / 0.699 0.911 / 0.750 0.732 0.898 / 0.690 0.906 / 0.739
Qwen2_5-32b 0.811 0.898 / 0.659 0.923 / 0.744 0.816 0.906 / 0.661 0.931 / 0.744
Yi-34b 0.652 0.777 / 0.575 0.831 / 0.666 0.652 0.726 / 0.513 0.781 / 0.572
Llama3_1-70b 0.885 0.849 / 0.493 0.946 / 0.679 0.886 0.841 / 0.485 0.941 / 0.666
Qwen2_5-72b 0.849 0.930 / 0.668 0.948 / 0.726 0.858 0.909 / 0.645 0.939 / 0.729
CMCQRD phi3_5 0.706 0.818 / 0.664 0.845 / 0.712 0.246 0.374 / 0.369 0.397 / 0.417
Llama3_2-3b 0.666 0.671 / 0.493 0.806 / 0.635 0.705 0.688 / 0.518 0.829 / 0.679
Qwen2_5-3b 0.662 0.791 / 0.622 0.816 / 0.659 0.744 0.864 / 0.673 0.884 / 0.704
Llama3_1-8b 0.752 0.583 / 0.415 0.800 / 0.602 0.781 0.650 / 0.454 0.828 / 0.628
Qwen2_5-14b 0.888 0.95 / 0.737 0.958 / 0.747 0.894 0.951 / 0.74 0.957 / 0.770
Qwen2_5-32b 0.908 0.954 / 0.728 0.970 / 0.785 0.912 0.949 / 0.743 0.969 / 0.813
Yi-34b 0.841 0.873 / 0.624 0.900 / 0.668 0.831 0.794 / 0.546 0.831 / 0.582
Llama3_1-70b 0.917 0.926 / 0.684 0.973 / 0.809 0.912 0.912 / 0.617 0.972 / 0.789
Qwen2_5-72b 0.914 0.962 / 0.737 0.971 / 0.766 0.912 0.955 / 0.739 0.973 / 0.793
Appendix B Large Language Models Used
@@
Table 6: Default precision and quantised LLMs used in the experiments. All models can be found at the HuggingFace Hub.
Model Name Default Precision 4-bit Quantisation
phi3_5-chat microsoft/Phi-3.5-mini-instruct unsloth/Phi-3.5-mini-instruct-bnb-4bit
Llama3_2-3b-chat meta-llama/Llama-3.2-3B-Instruct unsloth/Llama-3.2-3B-Instruct-bnb-4bit
Qwen2_5-3b-chat Qwen/Qwen2.5-3B-Instruct unsloth/Qwen2.5-3B-Instruct-bnb-4bit
Llama3_1-8b-chat meta-llama/Llama-3.1-8B-Instruct unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit
Qwen2_5-14b-chat Qwen/Qwen2.5-14B-Instruct unsloth/Qwen2.5-14B-Instruct-bnb-4bit
Qwen2_5-32b-chat Qwen/Qwen2.5-32B-Instruct unsloth/Qwen2.5-32B-Instruct-bnb-4bit
Yi-34b-chat 01-ai/Yi-34B-Chat unsloth/yi-34b-chat-bnb-4bit
Llama3_1-70b-chat meta-llama/Llama-3.1-70B-Instruct unsloth/Meta-Llama-3.1-70B-Instruct-bnb-4bit
Qwen2_5-72b-chat Qwen/Qwen2.5-72B-Instruct unsloth/Qwen2.5-72B-Instruct-bnb-4bit
Appendix C Results Overview
@@
Table 7 presents an overview of the results of all experiments, along with Table 8 which presents the performance our setup using the Bio_clinical Bert Embeddings [2]. The results using BERT embeddings are repeated to facilitate comparison. In our experiments, using the Bio_clinical Bert Embeddings consistently led to worse performance (higher RMSE) for all experiments.
Table 7: Root Mean Squared Error (RMSE, the lower the better) on the test set using different sets of features. Best overall results are highlighted in light green. All results are averaged over ten repetitions, with the standard deviation not exceeding 0.002.
Method Biopsychology USMLE CMCQRD CMCQRD_IRT
Baselines
Dummy Regressor 0.1667 0.3110 0.1833 9.7568
Best Literature Result [24] - 0.291 - 8.5
Linguistic Features Baseline [8] 0.1544 0.3147 0.1852 9.3335
Only Text
TF-IDF 0.1479 0.3092 0.1843 9.2531
BERT Embeddings 0.1498 0.3066 0.1843 8.9347
Only Uncertainty
Default 4-bit Default 4-bit Default 4-bit Default 4-bit
1st Token Probabilities 0.1473 0.1539 0.3041 0.2960 0.1770 0.1752 9.4161 9.5385
Choice Order Sensitivity 0.1543 0.1582 0.3155 0.3178 0.1788 0.1880 9.9787 9.8563
Both Uncertainty Features 0.1460 0.1538 0.3034 0.2968 0.1754 0.1761 9.3705 9.4899
Only Choice Similarity
General (all-MiniLM-L6-v2) 0.1895 0.3567 0.2226 12.2829
Medical (S-PubMedBert-MS-MARCO) 0.1883 0.3432 0.2146 11.4768
TF-IDF BERT TF-IDF BERT TF-IDF BERT TF-IDF BERT
Text and Uncertainty
First Token Probability (Default) 0.1361 0.1362 0.3037 0.2868 0.1668 0.1669 8.5661 8.2763
Choice Order Sensitivity (Default) 0.1378 0.1409 0.3100 0.2906 0.1658 0.1676 8.6663 8.4204
Both Uncertainty Features (Default) 0.1359 0.1365 0.3044 0.2864 0.1654 0.1669 8.5933 8.3009
First Token Probability (4-bit) 0.1365 0.1385 0.2851 0.2854 0.1680 0.1675 8.5392 8.3322
Choice Order Sensitivity (4-bit) 0.1309 0.1411 0.2951 0.2961 0.1671 0.1672 8.6218 8.3459
Both Uncertainty Features (4-bit) 0.1371 0.1388 0.2856 0.2846 0.1685 0.1670 8.5587 8.3728
Text, Both Uncertainties & Choice Similarity
General Sim. (Default) 0.1372 0.1367 0.2835 0.2862 0.1651 0.1669 8.5934 8.2956
Medical Sim. (Default) 0.1376 0.1361 0.2836 0.2853 0.1651 0.1664 8.5999 8.2325
Both Sim. (Default) 0.1378 0.1365 0.2847 0.2862 0.1653 0.1676 8.5694 8.2817
General Sim. (4-bit) 0.1386 0.1389 0.2850 0.2841 0.1674 0.1659 8.5696 8.3338
Medical Sim. (4-bit) 0.1378 0.1381 0.2856 0.2844 0.1676 0.1660 8.5204 8.2801
Both Sim. (4-bit) 0.1397 0.1412 0.2860 0.2850 0.1673 0.1659 8.5475 8.2820
Text, Both Uncertainties, Both Sim & Linguistic Features
Default 0.1393 0.1362 0.2817 0.2857 0.1652 0.1673 8.5490 8.6575
4-bit 0.1411 0.1410 0.2853 0.2844 0.1677 0.1660 8.4966 8.5907
@@
Table 8: Comparison in Root Mean Squared Error (RMSE, the lower the better) achieved using BERT or Bio_Clinical BERT embeddings. Experiments using the CMCQRD were not conducted, as the dataset does not cover biological/medical domain. All results are averaged over ten repetitions, with the standard deviation not exceeding 0.002.
Method Biopsychology USMLE
BERT Bio_Clinical BERT BERT Bio_Clinical BERT
Text and Uncertainty
First Token Probability (Default) 0.1362 0.1376 0.2868 0.2945
Choice Order Sensitivity (Default) 0.1409 0.1430 0.2906 0.3004
Both Uncertainty Features (Default) 0.1365 0.1387 0.2864 0.2952
First Token Probability (4-bit) 0.1385 0.1396 0.2854 0.2908
Choice Order Sensitivity (4-bit) 0.1411 0.1426 0.2961 0.3035
Both Uncertainty Features (4-bit) 0.1388 0.1392 0.2846 0.2909
Text, Both Uncertainties & Choice Similarity
General Sim. (Default) 0.1367 0.1380 0.2862 0.2951
Medical Sim. (Default) 0.1361 0.1381 0.2853 0.2936
Both Sim. (Default) 0.1365 0.1377 0.2862 0.2936
General Sim. (4-bit) 0.1389 0.1424 0.2841 0.2926
Medical Sim. (4-bit) 0.1381 0.1425 0.2844 0.2924
Both Sim. (4-bit) 0.1412 0.1422 0.2850 0.2922
Text, Both Uncertainties, Both Sim & Linguistic Features
Default 0.1362 0.138 0.2857 0.2939
4-bit 0.1410 0.1426 0.2844 0.2915